OpenStudy (anonymous):

if x approaching 3 from the left =22 and x approaching 3 from the right = 22 but the lim f(x) as x approaches 3 = 20. is the function continuous at x=3? why or why not

OpenStudy (anonymous):

? if the limit from the left is 22 and the limit from the right is 22, then the "limit' cannot be 20

OpenStudy (anonymous):

im given a piecewise

OpenStudy (anonymous):

f(x)= 2x^2 +4 when x<3 20 when x=3 28-2x when x>3

OpenStudy (anonymous):

and im asked to find the limit of x to 3+ x to 3- x to 3

OpenStudy (anonymous):

piecwise or not, if \[\lim_{x\rightarrow 3^-}f(x)=22=\lim_{x \rightarrow 3^+}f(x)\] then the limit is 22, not something else

OpenStudy (unklerhaukus):

the function is not continuous at the point x=3

OpenStudy (anonymous):

ah the limit is 22, but the function is not continuous at 3 because it is not the same as the value of the limit there

OpenStudy (anonymous):

why is that unklerhaukus

OpenStudy (anonymous):

for the function to be continuous, it must have the same value as the limit

OpenStudy (anonymous):

so since \[\lim_{x\rightarrow 3} f(x)=22\] but \[f(3)=20\] it is not continuous at 3

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

Satellite73 has already said this but The function is Only continuous at some point if the limit Exists And is Equal to the function at that point.

OpenStudy (anonymous):

ok can you help me with this one state the definition of a function f(x) which is continuous at x= a

OpenStudy (anonymous):

Yes, actually this (following) statement is pretty confusing "the lim f(x) as x approaches 3 = 20."

OpenStudy (anonymous):

i was wrong there

OpenStudy (anonymous):

A function is continuous at x=a if \[ \large \space \lim \limits_{x\rightarrow a^-}f(x)=f(a)=\lim \limits_{x \rightarrow a^+}f(x) \]

OpenStudy (unklerhaukus):

(assuming the limit exists of course)

OpenStudy (anonymous):

If \( \space \lim \limits_{x\rightarrow a^-}f(x)=\lim \limits_{x \rightarrow a^+}f(x) \) then the limit exists.